OpenStudy (anonymous):

What is the value of y? 6y2 + 5 = 221

OpenStudy (anonymous):

Is that 6y squared? (little 2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=6, -6

OpenStudy (anonymous):

it has this sign by it ±

OpenStudy (skullpatrol):

that is read "plus or minus"

OpenStudy (skullpatrol):

it means two answers that have opposite signs

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

6y^2+5-5=221-5 6y^2=216 6y^2/6=216/6 y^2=36 39 square root y=6 positive 6 (+6)

OpenStudy (skullpatrol):

http://en.wikipedia.org/wiki/Plus-minus_sign

OpenStudy (anonymous):

thx

OpenStudy (skullpatrol):

thx for asking :)

OpenStudy (anonymous):

Solve the equation. If necessary, round your answer to the nearest tenth. x2 = 256

OpenStudy (anonymous):

square root of 256 will be x

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

No problem :)